OpenStudy (anonymous):

For which value of x should the following expression be further simplified? sqrt(39x) A. x = 2 B. x= 6 C. x= 10 D. x = 11 i think its c??

OpenStudy (anonymous):

@StudyGirl14 could you help me again

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

B, actually

OpenStudy (studygurl14):

because \(\large\sqrt{39(6)}\rightarrow\sqrt{234}\rightarrow\sqrt{9\times26}\rightarrow 3\sqrt{26}\)

OpenStudy (anonymous):

thankyou!! :D:D:D